a rectangular 55.0 turns coil of wire of dimensions 50 cm by 30 cm. It carries a current of 1.5 A in the counterclockwise direction. It is mounted in the x-y plane. The magnetic field makes an angle of θ= 37.0° with the positive x-axis and magnitude 1.2 T. What is the magnitude of the torque (in Nm) acting on the coil.
Answer #1

|\tau| =BNIA Sin\theta

\theta is the angle between magnetic field and area vector ,

so torque=(1.2)*(55)*(1.5)*(0.5)*(0.3) sin(53)=11.86 Nm

Here angle between normal to area and magnetic fieldld is 53